# Hierarchical and Modular Network on Non-prehensile Manipulation in General Environments

### v1

@@ Abstract
For robots to operate in general environments like households, they must be able to perform non-prehensile manipulation actions such as toppling and rolling to manipulate ungraspable objects. However, prior works on non-prehensile manipulation cannot yet generalize across environments with diverse geometries. The main challenge lies in adapting to varying environmental constraints: within a cabinet, the robot must avoid walls and ceilings; to lift objects to the top of a step, the robot must account for the step’s pose and extent. While deep reinforcement learning (RL) has demonstrated impressive success in non-prehensile manipulation, accounting for such variability presents a challenge for the generalist policy, as it must learn diverse strategies for each new combination of constraints. To address this, we propose a modular architecture that uses different combinations of reusable modules based on task requirements. To capture the geometric variability in environments, we extend the contact-based object representation from CORN [14] to environment geometries, and propose a procedural algorithm for generating diverse environments to train our agent. Taken together, the resulting policy can zero-shot transfer to novel real-world environments despite training entirely within a simulator. We additionally release a simulation-based benchmark featuring nine digital twins of real-world scenes with 353 objects to facilitate non-prehensile manipulation research in realistic domains.
@@
11footnotetext: Equal Contributions.
@@ I Introduction
The key challenge here lies in training a policy capable of quickly adapting to the constraints imposed by the current scene. For instance, consider the scenarios in Figure 1: each domain not only presents the robot with a unique set of constraints, but these constraints also evolve throughout an episode (Figure 1e). This requires a high-frequency policy function that can model rapid changes in actions, even in response to minor changes in the state or goal. For example, in the scenario depicted in Figure 1e, as soon as the toy crab is positioned above the sink, the robot must quickly switch from a lifting skill to a translation skill. Unfortunately, neural networks struggle to learn such high-frequency functions, even in supervised learning setups [83, 3, 73], a phenomenon known as spectral bias.
@@ II-B Multi-task Neural Architectures
Representative context-adaptive architectures include conditional normalization [68, 66, 54, 5], hypernetworks [31, 36], and modular architectures [29, 94, 81]. In conditional normalization [68] (Figure 5b), $g_{\phi}$ takes the context variable $z$ as input and applies feature-wise scale and bias $\{\gamma,\beta\}=g_{\phi}(z)$ to the intermediate features of the base network as $y=\gamma\odot f_{\theta}(x)+\beta$ where $\odot$ denotes the element-wise multiplication. While effective, the expressivity of these architectures is limited to affine transforms, restricting its capacity [74]. In hypernetworks [31] (Figure 5c), $g_{\phi}$ affords broader expressivity, as it generates the entire set of base network parameters, $\theta=g_{\phi}(z)$, but suffer from poor training stability due to its large decision space over densely interacting parameters [60].
@@
However, prior works on modular architectures for control assume that $z$ is given by the user and is discrete, such as pre-defined task IDs [69, 81, 94, 33]. In contrast, to generalize to novel, real-world environments in our problem, we need to infer $z$ from sensory observations such as environment and object point clouds, and a goal. To overcome this, we design geometric encoders that map high-dimensional sensory observations to $z$.
Our architecture is most similar to Soft Modularization (SM) [94], with three key differences. Unlike SM, which predicts connectivities between all pairs of modules between neighboring layers, we directly predict module activations, which simplifies computation and reduces the output dimensions from $M^{2}$ connections to $M$ module activations. We additionally improve the computational efficiency by predicting all module activation weights in parallel, instead of predicting in series, conditioning on the preceding layer’s module activations. Lastly, we incorporate a gating mechanism [31] to enhance expressivity and boost policy performance.
@@ II-C Representation learning on point clouds
To accelerate RL training with high-dimensional sensory inputs, prior works use representational pre-training [95] to bootstrap RL agents. Different pretext tasks have been proposed for this purpose, such as point completion [88], orientation and category estimation [9, 35], or contrastive learning [93]. Inspired by the advances in natural language processing [19, 72] and image analysis [32], recent works adopt self-supervised learning (SSL) on patch-based transformers [97, 8, 65, 101, 1] for point cloud representation learning. These works reconstruct unseen geometric patches via either autoregressive prediction [8] or masking [97, 65, 101, 1] to learn rich geometric representations, achieving state-of-the-art results in shape classification and segmentation [8].
@@ III Methodology
We consider the non-prehensile manipulation problem, where a robot arm with a fixed base moves an object to a target pose in environments of general geometry, e.g., kitchen sinks, cabinets, and drawers (Figure 1). We model this problem as a Markov Decision Process (MDP), represented as a tuple $(S,A,P,r,\gamma)$ denoting state space $S$, action space $A$, state transition model $P(s_{t+1}|a_{t},s_{t})$, reward model $r(s_{t},a_{t},s_{t+1})$, and discount factor $\gamma$. Our objective is to obtain policy $\pi$ that maximizes the return $R_{t}=\mathbb{E}_{a_{t}\sim\pi(\cdot|s_{t})}[\sum\gamma^{t}r(s_{t},a_{t},s_{t+1})]$ via a sequence of non-prehensile actions. Details on our MDP design are in the Appendix A-D.
Figure 7 presents an overview of our framework. We leverage deep RL in a parallel GPU-based simulation [49] to train a modular policy (Section III-A) using the pre-trained point cloud representation (Section III-B) on procedurally generated domains (Section III-C). We distill the resulting policy for real-world deployment via teacher-student distillation (Appendix A-C). All pre-training, policy training, and distillation stages happen entirely in a simulation.
III-A HAMnet-based architecture
@@
The geometry encoder processes three types of point cloud inputs: the global scene cloud, capturing the overall geometry of the scene; the local scene cloud, detailing the nearby scene that surrounds the object; and the object point cloud, representing its surface geometry. Each cloud is patchfied, tokenized, and embedded by the pre-trained geometry encoder (Section III-B), yielding latent geometric embeddings $z^{(G)}_{\text{1:N}}$, $z^{(L)}_{\text{1:N}}$, $z^{(O)}_{\text{1:N}}$, for global, local, and object embeddings, respectively. Details on point cloud acquisition are in Appendix A-E.
@@ III-B Training UniCORN
III-B1 Pre-training data generation
To acquire data for pre-training, we generate a dataset containing pairs of objects represented as point clouds, and contact labels indicating the presence and location of contact. Using the objects from DexGraspNet dataset [89], we generate the data by (1) sampling near-contact object configurations, (2) creating the point clouds by sampling points from the surface of each object, and (3) labeling contact points based on whether they fall within the other object. To account for possible scale variations between geometries, we sample the point clouds at varying densities and scales during this process. For details on the data generation pipeline, see Appendix A-A.
III-B2 Network Architecture
@@
III-C Procedural domain and curriculum generation
@@
Our procedural pipeline, shown in Figure 9, comprises two main components: domain interior and exterior generation. The interior includes planar and interim plates arranged laterally and longitudinally, where planar plates form elevated surfaces and interim plates form sloped ramps. Their dimensions, elevations, and angles are randomly sampled to produce diverse topographies, yielding features like bumps, valleys, and steps (Figure 3). The exterior consists of walls and ceilings that impose accessibility constraints, where their presence, height, and ceiling type (nominal or tight) are randomly sampled. The proportion of ceiling types controls the difficulty of workspace accessibility, since the nominal ceiling is generated with sufficient clearance, whereas tight ceilings leave a narrow margin relative to the object’s height. Details on the procedural generation pipeline are in Appendix A-B.
@@ IV-A Overview
To evaluate our claims, we compare the performance of our proposed model (UniCORN-HAMnet) with the baselines summarized in Table II. These baselines explore alternative choices in network architecture or representation. UniCORN-Hyper uses a hypernetwork [42, 76] to predict base network parameters. UniCORN-SM is a variant of a modular architecture using Soft Modularization [94]. We include two variants of standard monolithic architectures, UniCORN-Mono and UniCORN-Transformer, respectively using an MLP and a transformer. Lastly, PointGPT-HAMnet replaces UniCORN with the PointGPT encoder [8] based on the pre-trained weights released by the authors. All architectures are configured to have a similar number of trainable parameters up to the architectural constraints. Additional details on the baselines are in Appendix A-F.
@@ IV-B Simulation experiment
To train our policy, we use a Franka Research 3 (FR3) arm manipulating a subset of 323 objects from DexGraspNet dataset [89] on the procedurally generated environments as in Section III-C. We train each baseline using PPO [77] with identical hyper-parameters, spanning 2 billion environment interactions across 1024 parallel environments in Isaac Gym [49]. Detailed hyperparameters for policy training are described in Appendix F. We consider two metrics: data efficiency and time efficiency.
@@
To support our claim on training efficiency, we consider the training progression plot in Figure 10. Overall, modular architectures (UniCORN-HAMNet and UniCORN-SM) achieve the best data efficiency, with the mean success rates of 75.6% and 70.9% after training. In contrast, monolithic architectures show lower performance, regardless of whether conditioning is given by concatenation (UniCORN-Mono, 49.8%) or self-attention (UniCORN-Transformer, 43.9%). UniCORN-Hyper performs best among non-modular architectures at 56.5%, indicating the adaptivity of the network expedites policy training. UniCORN-SM (70.9%) further improves over hypernetworks, as its modularity affords reuse of network modules and reduces the learning complexity by predicting sparse module activations rather than parameters of individual neurons. Lastly, UniCORN-HAMNet (75.6%) outperforms UniCORN-SM from the increased expressivity of the gating mechanism. To evaluate the representational efficacy of UniCORN, we also compare UniCORN-HAMNet to PointGPT-HAMNet, which underperforms UniCORN-HAMNet (66.0% vs. 75.6%) due to the overhead from spurious geometric details and increased embedding dimensions.
@@
Since all our training happens in simulation, training time is also an important factor. To evaluate the time-efficiency of each baseline, we measure the per-epoch training time in Figure 11. While the monolithic MLP architecture (UniCORN-Mono) is the fastest (1.12s) due to its simplicity, modular architectures (UniCORN-HAMNet and UniCORN-SM) follow closely at just 1.33s and 1.84s, which shows that determining module-level activation adds minimal overhead; between the two, UniCORN-HAMNet achieves faster training than UniCORN-SM from the streamlined prediction of module activations. In contrast, UniCORN-Hyper takes the longest (5.31s) due to the cost of predicting the full set of base network parameters. To contextualize the overhead, a standard transformer (UniCORN-Transformer) takes around 3.30s. We also compare with PointGPT-HAMNet, where the large encoder significantly lags training, averaging 4.40 s per epoch, about 3.3 times slower than UniCORN-HAMNet.
@@ IV-C Real world experiment
(a) Example of our scene layout in the cabinet domain.
(b) Our real-world objects, all unseen during policy training.
Figure 12: Our real-world experimental setup.
@@
To validate the real-world applicability and generalizability of our framework, we evaluate our policy in 9 real-world domains with novel everyday scenes and objects (Figure 1). We test two objects in each domain: one object (a toy bulldozer), shared across all domains, and one random object (Figure 12(b)), each with five trials at different initial and goal poses. We replaced the robot’s gripper to accommodate narrow environments, wrapped with a high-friction glove to reduce slipping. Details on the real-world setup are in Appendix A-H.
Table III shows the results of our policy across 9 real-world domains. Overall, our policy demonstrates 78.9% success rate, indicating that our framework facilitates the policy to transfer to diverse, unseen real-world environments, despite only training in a simulation. The main failure modes of our policy are, in decreasing order of frequency: torque limit violation (5.56%); policy deadlock (4.44%); dropping objects (4.44%); getting blocked by the environment (3.33%); and perception error (3.33%). Detailed descriptions of these failure modes are in Appendix A-H2.

### v2

@@ Abstract
For robots to operate in general environments like households, they must be able to perform non-prehensile manipulation actions such as toppling and rolling to manipulate ungraspable objects. However, prior works on non-prehensile manipulation cannot yet generalize across environments with diverse geometries. The main challenge lies in adapting to varying environmental constraints: within a cabinet, the robot must avoid walls and ceilings; to lift objects to the top of a step, the robot must account for the step’s pose and extent. While deep reinforcement learning (RL) has demonstrated impressive success in non-prehensile manipulation, accounting for such variability presents a challenge for the generalist policy, as it must learn diverse strategies for each new combination of constraints. To address this, we propose a modular architecture that uses different combinations of reusable modules based on task requirements. To capture the geometric variability in environments, we extend the contact-based object representation from CORN [17] to environment geometries, and propose a procedural algorithm for generating diverse environments to train our agent. Taken together, the resulting policy can zero-shot transfer to novel real-world environments despite training entirely within a simulator. We additionally release a simulation-based benchmark featuring nine digital twins of real-world scenes with 353 objects to facilitate non-prehensile manipulation research in realistic domains. Code, videos, and simulation benchmarks are available on the project website.
@@
11footnotetext: Equal Contribution.
@@ I Introduction
The key challenge here lies in training a policy that can adapt to the constraints imposed by the given scene. For instance, consider the scenarios in Figure 1: for each domain, the robot is presented with a unique set of constraints, which may also evolve during an episode (Figure 1e). This requires the policy to not only model multiple distinct behaviors, but also rapidly switch between behaviors in response to minor changes in state. For example, in the scenario depicted in Figure 1e, as soon as the toy crab is positioned above the sink, the robot must quickly switch from a lifting skill to a translation skill. However, standard networks struggle to learn such high-frequency functions, a phenomenon known as spectral bias [99, 4, 88].
@@ II-B Multi-task Neural Architectures
Representative context-adaptive architectures include conditional normalization [82, 80, 65, 7], hypernetworks [37, 42], and modular architectures [34, 112, 97]. In conditional normalization [82] (Figure 5b), $g_{\phi}$ takes the context variable $z$ as input and applies feature-wise scale and bias $\{\gamma,\beta\}=g_{\phi}(z)$ to the intermediate features of the base network as $y=\gamma\odot f_{\theta}(x)+\beta$ where $\odot$ denotes element-wise multiplication. While effective, the expressivity of these architectures is limited to affine transforms, restricting its capacity [89]. In hypernetworks [37] (Figure 5c), $g_{\phi}$ affords broader expressivity, as it generates the entire set of base network parameters, $\theta=g_{\phi}(z)$, but suffer from poor training stability due to its large decision space over densely interacting parameters [71].
@@
However, prior works on modular architectures for control assume that $z$ is given by the user and is discrete, such as pre-defined task IDs [84, 97, 112, 39]. In contrast, to generalize to novel, real-world environments in our problem, we need to infer $z$ from sensory observations such as environment and object point clouds. To overcome this, we design geometric encoders that map high-dimensional sensory observations to $z$.
Our architecture is most similar to Soft Modularization (SM) [112], with three key differences. Unlike SM, which predicts connectivities between all pairs of modules between neighboring layers, we directly predict module activations, which simplifies computation and reduces the output dimensions from $M^{2}$ connections to $M$ module activations. We additionally improve the computational efficiency by predicting all module activation weights in parallel, instead of predicting in series while conditioning on the preceding layer’s module activations. Lastly, we incorporate a gating mechanism [37] to enhance expressivity and boost policy performance.
@@ II-C Representation learning on point clouds
To accelerate RL training with high-dimensional sensory inputs, prior works use representational pre-training [114, 29] to bootstrap RL agents. Different pretext tasks have been proposed for this purpose, such as point completion [105], orientation and category estimation [12, 41], or contrastive learning [110]. Inspired by the advances in natural language processing [22, 87] and image analysis [38], recent works adopt self-supervised learning (SSL) on patch-based transformers [116, 11, 76, 120, 1] for point cloud representation learning. These works reconstruct unseen geometric patches via either autoregressive prediction [11] or masking [116, 76, 120, 1] to learn rich geometric representations, achieving state-of-the-art results in shape classification and segmentation [11].
@@ II-D Modularity in biological networks
II-E Skill discovery in RL
In skill discovery, Unsupervised RL (URL) aims to find reusable skills using task-agnostic objectives like state coverage [79, 51, 64, 83, 113, 8] or skill diversity [35, 26, 93, 47, 77]. However, without task-specific priors, such intrinsically motivated methods often fail to make meaningful interactions in high-dimensional domains without engineered bias [3, 102]. While METRA [78] was shown to scale to pixel-based tasks, it remains prone to degenerate behaviors, such as lying still in varied poses in humanoid control domains [102]. On the other hand, our framework learns task-specific skills grounded in the training domain, yielding interpretable skills that transfer to real-world robot manipulation.
II-F Mitigating Spectral Bias in RL
A central challenge in multi-task learning is enabling networks to adaptively switch functions under subtle shifts in its input. While our work focuses on the design of network architecture, one complementary line of work investigates improving input representations with high-frequency encodings, such as Learned Fourier Features (LFFs)[99], to enhance the network’s sensitivity to sharp transitions. While promising, the effectiveness of LFFs remains inconclusive in RL, with conflicting claims on the role of different frequency components [48, 111, 60]. Moreover, LFFs lead to an exponential increase in input dimensionality [6] and may degrade training stability from the increased variance, impairing generalization and robustness to noise [60]. These limitations suggest the need for an architecture-level solution, and not just the input representation. Still, one potentially promising future direction is to see whether high-frequency embeddings can complement the modular architecture by increasing network sensitivity to distinct contexts when determining module activations.
@@ III Methodology
We consider the non-prehensile manipulation problem, where a robot arm with a fixed base moves an object to a target pose in environments of general geometry, e.g., kitchen sinks, cabinets, and drawers (Figure 1). We model this problem as a Markov Decision Process (MDP), represented as a tuple $(S,A,P,r,\gamma)$ denoting state space $S$, action space $A$, state transition model $P(s_{t+1}|a_{t},s_{t})$, reward model $r(s_{t},a_{t},s_{t+1})$, and discount factor $\gamma$. Our objective is to obtain policy $\pi$ that maximizes the return $R_{t}=\mathbb{E}_{a_{t}\sim\pi(\cdot|s_{t})}[\sum\gamma^{t}r(s_{t},a_{t},s_{t+1})]$ via a sequence of non-prehensile actions.
Figure 6 presents an overview of our framework. We leverage deep RL in a parallel GPU-based simulation [58] to train a modular policy (Section III-B) using the pre-trained point cloud representation (Section III-C) on procedurally generated domains (Section III-D). We distill the resulting policy for real-world deployment via teacher-student distillation (Appendix -D). All pre-training, policy training, and distillation stages happen entirely in a simulation.
III-A MDP Design
Our state space $S$ consists of robot joint state $x^{q}_{t}$, end-effector pose $x_{t}^{EE}$, object geometry $G_{o}$, environment geometry $G_{e}$, and goal pose $T_{g}$. In the simulation, the agent additionally receives physics parameters $\nu$ and object state $x^{o}_{t}$. We represent all poses as 3D translation and 6D orientation to facilitate learning [123], and the goal is given as a relative pose from the current object pose. Object and scene geometries are given as surface-sampled point clouds.
The reward $r(s_{t},a_{t},s_{t+1})$ in our domain is defined as a sum of the task success reward $r_{s}$, goal-reaching reward $r_{r}$ and the contact-inducing reward $r_{c}$: $r=r_{s}+\lambda_{r}r_{r}+\lambda_{c}r_{c}$, where $\lambda_{r}$ and $\lambda_{c}$ are scaling coefficients for the respective rewards. Since $r_{s}=\mathds{1}_{suc}$ is sparsely given, we incorporate shaping rewards $r_{r}$ and $r_{c}$ as potential functions of the form $\gamma\phi(s^{\prime})-\phi(s)$ with the discount factor $\gamma\in[0,1)$, which preserves policy optimality [69]. Specifically, we have $\phi_{r}(s)=-\log{(c_{g}\cdot d_{o,g}(s)+1)}$ for $r_{r}$, and $\phi_{c}(s)=-\log{(c_{r}\cdot d_{h,o}(s)+1)}$ for $r_{c}$, where $c_{g},c_{r}\in\mathbb{R}$ are scaling coefficients for the distance-based potential functions; $d_{o,g}(s)$ is the relative distance between the current object and the goal pose, based on the bounding-box distance [2]; $d_{h,o}(s)$ is the hand-object distance between the object and the tip of the end-effector. Task success is achieved when the object’s pose is within 0.1m and 0.1 radians of the target pose. The episode terminates if (1) the object reaches the goal, (2) the object is dropped from the workspace, or (3) the episode reaches the timeout of 300 simulation steps. Table VII summarizes our MDP design, and details on reward coefficients are in Table VIII.
III-B HAMnet-based architecture
@@
The geometry encoder processes three types of point cloud inputs: the global scene cloud, capturing the overall geometry of the scene; the local scene cloud, detailing the nearby scene that surrounds the object; and the object point cloud, representing its surface geometry. Each cloud is patchfied, tokenized, and embedded by the pre-trained geometry encoder (Section III-C), yielding latent geometric embeddings $z^{(G)}_{\text{1:N}}$, $z^{(L)}_{\text{1:N}}$, $z^{(O)}_{\text{1:N}}$, for global, local, and object embeddings, respectively. Details on point cloud acquisition are in Appendix -F.
@@ III-C Training UniCORN
III-C1 Pre-training data generation
To acquire data for pre-training, we generate a dataset containing pairs of objects represented as point clouds, and contact labels indicating the presence and location of contact. Using the objects from DexGraspNet dataset [106], we generate the data by (1) sampling near-contact object configurations, (2) creating the point clouds by sampling points from the surface of each object, and (3) labeling contact points based on whether they fall within the other object. To account for possible scale variations between geometries, we sample the point clouds at varying densities and scales during this process. For details on the data generation pipeline, see Appendix -B.
III-C2 Network Architecture
@@
III-D Procedural domain and curriculum generation
@@
Our procedural pipeline, shown in Figure 9, comprises two main components: domain interior and exterior generation. The interior includes planar and interim plates arranged laterally and longitudinally, where planar plates form elevated surfaces and interim plates form sloped ramps. Their dimensions, elevations, and angles are randomly sampled to produce diverse topographies, yielding features like bumps, valleys, and steps (Figure 3). The exterior consists of walls and ceilings that impose accessibility constraints, where their presence, height, and ceiling type (nominal or tight) are randomly sampled. The proportion of ceiling types controls the difficulty of workspace accessibility, since the nominal ceiling is generated with sufficient clearance, whereas tight ceilings leave a narrow margin relative to the object’s height. Details on the procedural generation pipeline are in Appendix -C.
@@ IV-A Overview
E2E-HAMnet HAMnet End-to-end
To evaluate our claims, we compare the performance of our proposed model (UniCORN-HAMnet) with the baselines summarized in Table II. These baselines explore alternative choices in network architecture or representation. UniCORN-Hyper uses a hypernetwork [50, 91] to predict base network parameters. UniCORN-SM is a variant of a modular architecture using Soft Modularization [112]. We include two variants of standard monolithic architectures, UniCORN-Mono and UniCORN-Transformer, respectively using an MLP and a transformer. PointGPT-HAMnet and E2E-HAMnet considers alternative choices in the representation, where the former replaces the pretrained UniCORN with a PointGPT encoder [11], while the latter jointly trains the representation model end-to-end. All architectures are configured to have a similar number of trainable parameters up to the architectural constraints. Additional details on the baselines are in Appendix -G.
@@ IV-B Simulation experiment
To train our policy, we use a Franka Research 3 (FR3) arm manipulating a subset of 323 objects from DexGraspNet dataset [106] on the procedurally generated environments as in Section III-D. We train each baseline using PPO [92] with identical hyper-parameters, spanning 2 billion environment interactions across 1024 parallel environments in Isaac Gym [58]. Detailed hyperparameters for policy training are described in Appendix -G. We consider two metrics: data efficiency and time efficiency.
@@
To support our claim on training efficiency, we consider the training progression plot in Figure 10. Overall, modular architectures (UniCORN-HAMNet and UniCORN-SM) achieve the best data efficiency, with the mean success rates of 75.6% and 70.9% after training. In contrast, monolithic architectures show lower performance, regardless of whether conditioning is given by concatenation (UniCORN-Mono, 49.8%) or self-attention (UniCORN-Transformer, 43.9%). UniCORN-Hyper performs best among non-modular architectures at 56.5%, indicating the adaptivity of the network expedites policy training. UniCORN-SM (70.9%) further improves over hypernetworks, as its modularity affords reuse of network modules and reduces the learning complexity by predicting sparse module activations rather than parameters of individual neurons. Lastly, UniCORN-HAMNet (75.6%) outperforms UniCORN-SM from the increased expressivity of the gating mechanism. To evaluate the representational efficacy of UniCORN, we also compare UniCORN-HAMNet to E2E-HAMNet and PointGPT-HAMNet. End-to-end training (E2E-HAMNet) degrades performance (47.9%), emphasizing the utility of pre-training; while PointGPT-HAMNet performs better (66.0%), it still underperforms UniCORN-HAMNet due to the overhead from spurious geometric details and increased embedding dimensions.
@@
Since all our training happens in simulation, training time is also an important factor. To evaluate the time-efficiency of each baseline, we measure the per-epoch training time in Figure 11. While the monolithic MLP architecture (UniCORN-Mono) is the fastest (1.12s) due to its simplicity, modular architectures (UniCORN-HAMNet and UniCORN-SM) follow closely at just 1.33s and 1.84s, which shows that determining module-level activation adds minimal overhead; between the two, UniCORN-HAMNet achieves faster training than UniCORN-SM from the streamlined prediction of module activations. In contrast, UniCORN-Hyper takes the longest (5.31s) due to the cost of predicting the full set of base network parameters. To contextualize the overhead, a standard transformer (UniCORN-Transformer) takes around 3.30s. We also compare with representational baselines: PointGPT-HAMNet and E2E-HAMNet. The large encoder of PointGPT-HAMNet significantly lags training, averaging 4.40s per epoch, about 3.3 times slower than UniCORN-HAMNet. While E2E-HAMNet uses the same encoder architecture as UniCORN-HAMNet, it suffers from 2.9$\times$ slower training due to the overhead of co-training the representation model.
@@ IV-C Real world experiment
To validate the real-world applicability and generalizability of our framework, we evaluate our policy in 9 real-world domains with novel everyday scenes and objects (Figure 1). We test two objects in each domain: one object (a toy bulldozer), shared across all domains, and one random object (Figure 12(b)), each with five trials at different initial and goal poses.
For each scene, we mount four RealSense D435 cameras to observe the point clouds from multiple viewpoints, ensuring sufficient visibility of the object during execution (Figure 12(a)). To distinguish the object cloud from the environment cloud, we use SAM [46] to designate the initial object segmentation mask and utilize Cutie [13] to track the object during manipulation. We use FoundationPose [108] to estimate the object’s relative pose from the goal pose, using the view with the best visibility of the object (largest object segmentation mask) among the four cameras. We generate the environment point cloud by combining and filtering the point clouds from the depth cameras. We replaced the robot’s gripper to accommodate narrow environments, wrapped with a high-friction glove to reduce slipping. Further details on the real-world setup are in Appendix -I.
Table III shows the results of our policy across 9 real-world domains. Overall, our policy demonstrates 78.9% success rate, indicating that our framework facilitates the policy to transfer to diverse, unseen real-world environments, despite only training in a simulation. The main failure modes of our policy are, in decreasing order of frequency: torque limit violation (5.56%); policy deadlock (4.44%); dropping objects (4.44%); getting blocked by the environment (3.33%); and perception error (3.33%). Detailed descriptions of these failure modes are in Appendix -I2.
(a) Example of our scene layout in the cabinet domain.
(b) Our real-world objects, all unseen during policy training.
Figure 12: Our real-world experimental setup.